OpenStudy (anonymous):

When the tax on cigarettes was increased by 5%, the price of a certain pack became $5.60 . What was the original price?

OpenStudy (anonymous):

Thanks for coming ganeshie

ganeshie8 (ganeshie8):

Let the original price be \(x\)

OpenStudy (anonymous):

Ok

ganeshie8 (ganeshie8):

Increasing \(x\) by \(5\%\) gives \($5.60\) : \[x + 0.05x = 5.60\] can you solve \(x\) from above ?

OpenStudy (anonymous):

No I think it would go like this x + 5. x. = 5.0 100.

OpenStudy (anonymous):

Am I right ?

ganeshie8 (ganeshie8):

That doesn't look correct

ganeshie8 (ganeshie8):

\(x + 0.05x = 5.60 \) \(1.05x = 5.60 \) \(x = \dfrac{5.60}{1.05}\) simplify

OpenStudy (anonymous):

5.33

OpenStudy (anonymous):

How did u get 0.05x

OpenStudy (anonymous):

Ganeshie

OpenStudy (astrophysics):

5% = 0.05

OpenStudy (mathmale):

@American.Eagle: Would you please ask clarifying questions, instead of just typing someone's name (ganeshie8). Explain what you know. Try to put into words whatever it is that you don't understand.